OpenStudy (anonymous):

consider the vectors u=(2,0) and v=(0,1) , 1.determine cos theta,where theta is the angle between u and v 2.determine the area of the parallelogram by u and v

OpenStudy (dumbcow):

oh sorry i didn't see this question \[\cos \theta = \frac{u*v}{|u||v|}\] u*v = (2*0) +(0*1) = 0 \[\rightarrow \cos \theta = 0\]

OpenStudy (dumbcow):